Find the area of the region that lies inside the circle r = 3sin"\theta" and outside the cardioid r = 1+sin"\theta"


1
Expert's answer
2022-05-08T11:42:56-0400


The cardioid intersects with the circle


"3\sin\theta=1+\sin \theta"

"\sin \theta=\dfrac{1}{2}"

The cardioid intersects with the circle at "(\dfrac{3}{2},\dfrac{\pi}{6}),(\dfrac{3}{2},\dfrac{5\pi}{6})" and the pole.

The area of interest has been shaded above.

To find the area of a polar curve, we use


"A=\dfrac{1}{2}\displaystyle\int_{\pi/6}^{5\pi/6}((3\sin \theta)^2-(1+\sin \theta)^2)d\theta"

"=\dfrac{1}{2}\displaystyle\int_{\pi/6}^{5\pi/6}(8\sin^2 \theta-2\sin\theta-1)d\theta"

"=\dfrac{1}{2}\displaystyle\int_{\pi/6}^{5\pi/6}(3-4\cos2 \theta-2\sin\theta)d\theta"

"=\dfrac{1}{2}[3\theta-2\sin2\theta+2\cos \theta]\begin{matrix}
 5\pi/6\\
 \pi/6
\end{matrix}"

"=\dfrac{1}{2}(\dfrac{5\pi}{2}+\sqrt{3}-\sqrt{3}-\dfrac{\pi}{2}+\sqrt{3}-\sqrt{3})"

"=\pi"

"\pi" square units.